what is the pOH of the solution when 1 M NaCN and 0.5 M HCl mixed in equal volume?

Grade:11

1 Answers

ARYAN BHUSHAN
6 years ago
1V-0.5V=NV
N=0.5 i.e 0.5 M or N(n=1) H+ ion are present in the solution  after formation of NaCl salt.
pH=-log(0.5)=0.3
therefore,pOH=14-pH=13.7